# EffLiFe: Efficient Light Field Generation via Hierarchical Sparse Gradient Descent

### v1

@@ 3.5 Implementation Details
Configuration Soft3D[39] ${}^{+}$[18] DeepView[1] EffLiFe EffLiFe-4I
Small baseline 0.9260 0.8884 0.9541 0.9393 0.9576
Medium baseline 0.9300 0.8874 0.9544 0.9272 0.9461
Large baseline 0.9312 0.8673 0.9485 0.9099 0.9261
Generation Time, sec - - $\sim$20.00 0.09 0.10
@@ 4.1 Baselines, datasets and metrics
Baselines. In order to evaluate the visual performance and real-time efficiency of our method, we compare it with four generalizable novel view synthesis methods: IBRNet[12], MVSNeRF[13], and ENeRF[24]. Additionally, we compare our method with five light field generation methods, namely Stereo Magnification (StereoMag)[18], LLFF[19], Soft3D[39], MLI[30], and DeepView[1]. Among the baseline methods, IBRNet, MVSNeRF and ENeRF conduct extra finetuning experiments to improve performance. However, finetuning takes so much time that it is incompatible with real-time light field generation, thus we only compare the results without finetuning.
@@
Datasets. For training and evaluation, we select 4 datasets: Spaces[1], Real Forward-Facing[19], SWORD[34] and Shiny[38] dataset. To compare results on the Real Forward-Facing evaluation set, comprising 8 scenes, and the SWORD evaluation set, comprising 25 scenes (selected from the total evaluation scenes) and Shiny evaluation set, comprising 8 scenes, we trained our method using a combined training set in the same manner as IBRNet[12]. This combined training set consists of 90 scenes from the Spaces training set and 35 scenes from the Real Forward-Facing training set. To compare results on the Spaces evaluation set, which consists of 10 scenes with three baseline settings of large, medium, and small, we trained our method using the same training dataset as DeepView, comprising 90 scenes from the Spaces training set.
@@
Metrics. We compared the visual performance of the proposed method using standard metrics, structural similarity (SSIM), peak signal-to-noise ratio (PSNR), and perceptual similarity (LPIPs). We evaluated the temporal efficiency of all methods based on the time required to generate a single representation (RGB images for IBRNet[12], MVSNeRF[13] and ENeRF[24], MPIs for Soft3D[39], StereoMag[18], DeepView[1] and our method, multi-layered meshes for MLI[30]). Furthermore, given that light field generation methods yield light field representations suitable for offline rendering, we compared the offline rendering speed of light field generation methods, also in FPS. (We use the OPENGL renderer provided by LLFF[19] evaluate the offline rendering speed of our method.) To assess the viability of each method for 3D display applications, we evaluated their 3D display efficiency (measured in FPS), which we define as the reciprocal of the time required to provide sufficient light field information to support light field rendering on a 3D display. Specifically, a 3D display (e.g., a looking glass) requires $n$ views ($n$ is set to 18 in Tab. I) at different viewing angles to reproduce a light field, providing a continuous, forward-facing viewing experience. Novel view synthesis methods necessitate $n$ forward processes to generate $n$ new views. In contrast, light field generation methods, which create light field representations, require only a single forward process to generate $n$ novel views because the built-up representations enable really fast novel view rendering in the forward-facing viewing range.
@@ 4.2 Experiment Configurations
We compare our method with state-of-the-art generalizable novel view synthesis methods and light field generation methods on public datasets including SWORD[34], Real Forward-Facing[19], Shiny[38] and Spaces [1].
SWORD, Real Forward-Facing and Shiny. To compare results on a sparse view setting, we set the number of input views to 3 for all methods. The evaluation image resolution for three datasets is $378\times 512$, and the number of MPI planes for our method is set to 40. To compare with MLI[30], we choose their best-performance model SIMPLI-8L that uses 8 layered meshes for representations. We train two versions of our models, EffLiFe and EffLiFe-4I, containing 3 and 4 update iterations respectively (refer to Sec. 4.5 for details about EffLiFe-4I). In order to assess the interpolation and extrapolation capabilities, we evaluate all views from the three datasets. For each view of a given scene, we select the nearest available 3 views as input. The evaluation results for a single scene are calculated by averaging the results across all views. Likewise, the final evaluation results for a dataset are computed by averaging the results across all scenes.
@@ 4.3 Quantitative Results
Compared with offline methods, our model EffLiFe-4I outperforms offline methods in terms of visual metrics, achieving top-1 or top-2 rankings on three datasets. And our default model EffLiFe strikes a balance between the rendering quality and efficiency. Our models, on average, achieve a 100×100$\times$100 × faster generation speed compared to other offline methods. Compared with the online method ENeRF[24], our models are a little bit slower in generation time; however, our method exhibits superior visual metrics on three datasets. When considering the metric of 3D Display Efficiency, our method stands out as the only one capable of supporting real-time light field video display. This is due to its high temporal efficiency in both light field generation and offline rendering speed.
Tab. II lists the quantitative results on the Spaces evaluation set, where we can see that DeepView generates the highest-quality rendering results overall, followed by our method and Soft3D[39]. However, in terms of time efficiency, our method is about 200 times faster than DeepView. The Generation Time of Soft3D and ${}^{+}$ are left as ”-” because Soft3D is not open-sourced and ${}^{+}$ is a modified version by the authors of DeepView, so that we can not evaluate their time efficiency accurately. To summarize, we have the following observations:
  • Compared with offline novel view synthesis methods[12, 13] and offline light field generation methods[18, 39, 1, 30], our approach achieves around 100×100$\times$100 × speedup ratio with superior or comparable visual performance.
@@ 4.4 Qualitative Results
Among all the evaluation metrics, our proposed approach achieves comparable or better visual quality with offline approaches, and performs superior to online approaches. More qualitative comparison results on Shiny[30] and IBRNet collected[12] are shown in Fig. 9. Our method produces clear results that accurately captures view-dependent effects, albeit with slightly less sharpness compared to SIMPLI-8L[30]. This trade-off is made in exchange for reduced time spent constructing the light field representations.
@@ 4.5 Ablation Study
Number of iterations and planes, and backbone network. Tab. III demonstrates that if we only consider image quality, using as many MPI planes and iterations as possible would be ideal. However, to strike a balance between rendering quality and time efficiency, our default model configuration consists of only 3 iterations and 40 MPI planes. In the doubled conv layers setting, we doubled the number of convolution layers and increased the number of channels to improve visual metrics. While this change decreases the rendering FPS, it further enhances the rendering quality. Nevertheless, a more efficient and robust backbone network could be implemented to improve both rendering speed and quality.
@@
Sparsity Loss. Fig. 10 visually compares the differences in the rendered images and individual MPI planes from experiments conducted with and without the sparsity loss. Compared to the model trained with sparsity loss, the model trained without it generates rendered images that display more artifacts, particularly around sharp edges. A comparison of individual MPI planes shows that those produced by the model with sparsity loss contain clearer surfaces. This evidence suggests that sparsity loss aids the network in minimizing scene content layout ambiguity, resulting in cleaner and sharper images.
Figure 11: The qualitative effects of the occlusion refinement module. The left column is the rendered image by EffLiFe. The ”with occ” column stands for the results rendered using the occlusion refinement module, and the ”w/o occ” column stands for the results without using it.
Occlusion Refinement. In Fig. 11, we compare the results between experiments trained with and without the iterative occlusion refinement module. The occlusion-aware iterative refinement module markedly enhances the visibility of background content adjacent to the foreground edges. However, this module is deactivated in the model as the occlusion in the evaluation datasets is generally mild, and its activation could decrease the model’s frames per second (FPS).
@@
w/o sparsity loss 31.73 24.98 0.81 0.20
w occlusion refinement 17.46 25.24 0.84 0.16
$m=4$(4 iters) 28.24 25.97 0.88 0.15
$D=16$(fewer planes) 47.35 23.38 0.79 0.22
$D=64$(more planes) 20.79 25.87 0.88 0.16
doubled conv layers 20.69 25.81 0.88 0.17
default model 31.73 25.42 0.85 0.16
@@ 5 Conclusion
This paper presents a novel method for efficiently generating light fields using hierarchical sparse gradient descent. Our proposed method achieves a resolution of $378\times 512$ for light field representations (MPIs) at a frame rate of approximately 30 FPS, from which novel views can be rendered at around 700 FPS. The hierarchical sparse gradient descent module of our network focuses on important MPI gradients, resulting in significant improvements in time efficiency without compromising rendering quality. And the occlusion-aware iterative refinement module helps eliminate severe occlusion, further enhancing rendering quality to state-of-the-art levels. Our method has the potential to deliver high-quality and real-time light field videos for XR and Naked Eye 3D display devices.
Limitations. Our approach has a few limitations. Firstly, our current implementation can only be trained with a fixed number of views, and the order of input views may slightly affect the final rendering quality. Secondly, our model is unable to produce relatively good results at the borders of the image where not all source views are overlapped, shown in Fig. 12.
@@
Future work. Although the proposed approach is able to generate light fields at around 30 FPS of resolution $378\times 512$, its efficiency could still be improved with customized CUDA kernels, which is not emphasized in this paper. Furthermore, a robust and adaptive view-aggregating module is required to support an arbitrary number of input views. This enhancement will facilitate the use of our model for various camera configurations.

### v3

@@ 4 Implementation Details
Configuration Soft3D[49] ${}^{+}$[21] DeepView[1] RealLiFe RealLiFe-2I
Small baseline 0.9260 0.8884 0.9541 0.9396 0.9286
Medium baseline 0.9300 0.8874 0.9544 0.9317 0.9159
Large baseline 0.9312 0.8673 0.9485 0.9180 0.9052
Generation Time, sec - - $\sim$20.00 0.044 0.037
@@ 5.1 Baselines, datasets and metrics
Baselines. In order to evaluate the visual performance and real-time efficiency of our method, we compare it with four generalizable novel view synthesis methods: IBRNet[5], MVSNeRF[6], and ENeRF[7]. Additionally, we compare our method with five light field reconstruction methods, namely Stereo Magnification (StereoMag)[21], Soft3D[49], MLI[8], and DeepView[1]. Among the baseline methods, IBRNet, MVSNeRF and ENeRF conduct extra finetuning experiments to improve performance. However, finetuning takes so much time that it is incompatible with real-time light field reconstruction, thus we only compare the results without finetuning. LLFF[9] is excluded as a baseline method because the number of input views for the released model is set fixed to 5. Additionally, LLFF fuses multiple MPIs to render novel views, leveraging information of up to 10-20 input views, making it a method with dense view inputs. However, it is essential to acknowledge that LLFF[9] remains a robust approach for generating high-quality light fields offline.
@@
Datasets. For training and evaluation, we select 4 datasets: Spaces[1], Real Forward-Facing[9], SWORD[40] and Shiny[48]. To compare results on the Real Forward-Facing evaluation set, comprising 8 scenes, the SWORD evaluation set, comprising 25 scenes (selected from the total evaluation scenes) and Shiny evaluation set, comprising 8 scenes, we trained our method using a combined training set in the same manner as IBRNet[5]. This combined training set consists of 90 scenes from the Spaces training set and 35 scenes from the Real Forward-Facing training set. To compare results on the Spaces evaluation set, which consists of 10 scenes with three settings of large, medium, and small baselines, we trained our method using the same training dataset as DeepView[1], comprising 90 scenes from the Spaces training set.
@@
Metrics. We compared the visual performance of the proposed method using standard metrics, structural similarity (SSIM), peak signal-to-noise ratio (PSNR), and perceptual similarity (LPIPs). We evaluated the temporal efficiency of all methods based on Generation Time, the time required to generate a single representation (RGB images for IBRNet[5], MVSNeRF[6] and ENeRF[7], MPIs for Soft3D[49], StereoMag[21], DeepView[1] and our method, multi-layered meshes for MLI[8]). Furthermore, given that light field reconstruction methods yield light field representations suitable for offline rendering, we compared the offline rendering speed of light field reconstruction methods, also in FPS. (We use the OpenGL renderer provided by LLFF[9] to evaluate the offline rendering speed.) To assess the viability of each method for 3D display applications, we evaluated their 3D display efficiency (measured in FPS), which we define as the reciprocal of the time required to provide sufficient light field information to support light field rendering on a 3D display. Specifically, a 3D display (e.g., a looking glass) requires $n$ views ($n$ is set to 18 in Tab. I) at different viewing angles to reproduce a light field, providing a continuous, forward-facing viewing experience. Novel view synthesis methods necessitate $n$ forward processes to generate $n$ new views. In contrast, light field reconstruction methods require only a single forward process to generate $n$ novel views because the built-up representations enable fast novel view rendering in the forward-facing viewing range.
@@ 5.2 Experiment Configurations
We compare our method with state-of-the-art generalizable novel view synthesis methods and light field reconstruction methods on public datasets including SWORD[40], Real Forward-Facing[9], Shiny[48] and Spaces [1].
SWORD, Real Forward-Facing and Shiny. To compare results on a sparse view setting, we set the number of input views to 3 for all methods. The evaluation image resolution for three datasets is $378\times 512$, and the number of MPI planes for our method is set to 40. To compare with MLI[8], we choose their best-performance model SIMPLI-8L that uses 8 layered meshes for representations. We trained two versions of our models, RealLiFe and RealLiFe-2I. RealLiFe has 3 iterations of learned gradient descent, while RealLiFe-2I has 2 iterations, directly upsamping the 1/4-scale MPI to the original scale. In order to assess the interpolation and extrapolation capabilities, we evaluate all views from the three datasets. For each view of a given scene, we select the nearest available 3 views as input. The evaluation results for a single scene are calculated by averaging the results across all views. Likewise, the final evaluation results for a dataset are computed by averaging the results across all scenes.
@@ 5.3 Quantitative Results
Compared with offline methods, our model RealLiFe and RealLiFe-2I stands out in terms of visual metrics, achieving top-1 or top-2 rankings on three datasets. And our default model RealLiFe strikes a balance between the rendering quality and efficiency. Our models, on average, achieve a 100×100$\times$100 × faster generation speed compared to other offline methods. Compared with the online method. Our models are a little bit faster than ENeRF[7] in generation time; however, our method exhibits superior visual metrics on three datasets. When considering the metric of 3D Display Efficiency, our method stands out as the only one capable of supporting real-time light field video display. This is due to its high temporal efficiency in both light field reconstruction and offline rendering speed.
Tab. II lists the quantitative results on the Spaces evaluation set, where we can see that DeepView generates the highest-quality rendering results overall, followed by our method and Soft3D[49]. The decrease in rendering quality may be attributed to the fact that our method depends on network connections to propagate visibility information. As we increase the baseline, there is a corresponding increase in the number of network connections required to effectively propagate visibility, as noted in [1]. However, in terms of time efficiency, our method is about 400 times faster than DeepView. The Generation Time of Soft3D and ${}^{+}$ are left as ”-” because Soft3D is not open-sourced and ${}^{+}$ is a modified version by the authors of DeepView, so that we can not evaluate their time efficiency accurately. To summarize, we have the following observations:
  • Compared with offline novel view synthesis methods[5, 6] and offline light field reconstruction methods[21, 49, 1, 8], our approach achieves around 100×100$\times$100 × speedup ratio with superior or comparable visual performance.
@@ 5.4 Qualitative Results
Among all the evaluation metrics, our approach achieves comparable or better visual quality with offline approaches, and performs superior to online approaches. More qualitative comparison results on Shiny[8] and IBRNet collected[5] are shown in Fig. 10. Our method produces clear results that accurately captures view-dependent effects, albeit with slightly less sharpness compared to SIMPLI-8L[8]. This trade-off is made in exchange for reduced time spent constructing the light field representations.
@@ 5.5 Ablation Study
Sparsification factor $k$. It is evident that increasing the value of $k$ results in enhanced rendering performance but at the expense of slower generation speed. Hence, it is crucial to identify an appropriate $k$ that strikes a balance between rendering quality and efficiency. Notably, as $k$ surpasses the value of 7, the demand for GPU memory during the training process exceeds the capacity of an RTX 3090 GPU. To circumvent this limitation, models with $k$ exceeding 7 have to be trained and inferred in smaller patches, which may inevitably downgrades the rendering quality. So we conducted ablation on $k$ with values 3, 5 and 7 in Tab. III. The results confirm that an increasing number of $k$ results in marginal increase in visual quality but obvious decrease in generation speed.
Number of iterations. Comparing the configurations of the default model with that of L=2 (2 iterations) in Tab. III, it is evident that employing additional iterations of learned gradient descent yields only a marginal enhancement in rendering quality. However, this refinement results in a notable decrease of approximately 8 FPS in temporal efficiency. Therefore, to strike a balance between rendering quality and efficiency, employing a 2-iteration model suffices. This efficiency trade-off stems from the diminishing returns of increased iterations, where the incremental improvement in rendering quality does not adequately offset the cost in terms of temporal performance, hence motivating the preference for the 2-iteration model.
Number of planes. When comparing the configurations of the default model, D=16 (fewer planes), and D=64 (fewer planes) in Tab. III, it becomes evident that employing too few planes results in a substantial deterioration in rendering quality, albeit compensating for expedited generation speed. Conversely, an increase in the number of planes yields only marginal enhancements in rendering quality while significantly diminishing temporal efficiency. This observation suggests the presence of a potential saturation point for the number of MPI planes. Beyond this point, the incremental addition of planes may not yield linear improvements in performance.
@@
Sparsity Loss. The default model demonstrates superior rendering quality compared to the w/o sparsity loss in Tab. III. And Fig. 11 visually compares the differences in the rendered images and individual MPI planes from experiments conducted with and without the sparsity loss. Compared to the model trained with sparsity loss, the model trained without it generates rendered images that display more artifacts, particularly around sharp edges. A comparison of individual MPI planes shows that those produced by the model with sparsity loss contain clearer surfaces. This evidence suggests that sparsity loss aids the network in minimizing scene content layout ambiguity, resulting in cleaner and sharper images.
Backbone network. Unlike the most straightforward way to process the sparse MPI gradients with an MLP, we employ a 3D CNN for better performance as discussed in Sec. 3.3. We can see from Tab. III that the configuration of MLP backbone network results in a reduction in rendering quality when compared to the default model. The diminished quality is attributed to the limited receptive field of an MLP. Given that the default 3D CNN structure aims to strike a balance between rendering quality and efficiency, further exploration of a more robust and efficient architecture is warranted to enhance performance.
Spasification criterion. In Section 3.3, we referenced an MVS[24, 25, 26, 27, 28, 29]-style gradient sparsification approach that involves selecting the layer with the highest alpha gradient along the depth axis as well as its adjacent $(k-1)/2$ layers to form a sparsified volume of $k$ layers of gradient data. This strategy, however, only takes geometry into consideration and may not adequately capture certain light field intricacies, such as semi-transparent objects and thin structures. An overall comparison in Table III reveals that the MVS-sampling strategy exhibits lower rendering quality when compared with the default model.
@@
$k=3$ 43.29 24.95 0.83 0.17
$k=7$ 32.79 25.51 0.84 0.15
$D=16$(fewer planes) 54.64 24.39 0.79 0.19
$D=64$(more planes) 28.57 25.48 0.84 0.15
$L=2$(2 iterartions) 45.45 25.33 0.84 0.16
w/o sparsity loss 31.73 25.20 0.81 0.16
MLP 48.54 24.86 0.82 0.17
MVS-sampling 37.04 24.34 0.79 0.19
default model 37.04 25.46 0.84 0.15
@@ 6 Conclusion
This paper presents a novel method for efficiently reconstructing light fields using hierarchical sparse gradient descent. Our proposed method achieves a resolution of $378\times 512$ for light field representations (MPIs) at a frame rate of above 35 FPS, from which novel views can be rendered at around 700 FPS. The hierarchical sparse gradient descent module of our network focuses on scene-aligned sparse MPI gradients, resulting in significant improvements in temporal efficiency without compromising rendering quality. Our method has the potential to deliver high-quality and real-time light field videos for XR and Naked Eye 3D display devices.
Limitations. There are some limitations of our approach. Firstly, our current implementation can only be trained with a fixed number of views, and the order of input views may slightly affect the final rendering quality. Secondly, our model is unable to produce relatively good results at the borders of the image where not all source views are overlapped, shown in Fig. 12.
@@
Future work. Although the proposed approach is able to generate light fields at above 35 FPS of resolution $378\times 512$, its efficiency could still be improved with customized CUDA kernels, which is not emphasized in this paper. Furthermore, a robust and adaptive view-aggregating module is required to support an arbitrary number of input views. This enhancement will facilitate the use of our model for various camera configurations.